73-x=85-2x. What does x equal?
Answer by lenny460(1073)

73 - x = 85 - 2x. What does x equal ?
Add x to each side of the equation.
73 - x + x = 85 - 2x + x
-x + x = 0
-2x + x = -1x (or -x)
Therefore:
73 = 85 - x
Subtract 85 from each side of the equation.
73 - 85 = 85 - x - 85
73 - 85 = -12
85 - 85 = 0
Therefore:
-12 = -x or
-1x = -12
Divide each side by -1
-1x/-1 = -12/-1
x = 12


Answer:
x = 12




Lennox Obuong
Algebra Tutor
Nairobi, Kenya